the carnival is in town. We paid 33 dollars for tickets. There were 5 children and 4 adults. the adult tickets were 1.50 more than the children tickets. write an equation to find the price of the adult tickets then solve

Answer by richwmiller(17219)

4*a+5*c=33
a=1.5+c
4*(1.5+c)+5*c=33
6.0+4c+5*c=33
9*c=27.0
c=3.0
a=1.5+c
a=$4.50 for 4 adults and c=$3.00 for 5 children
check
a=1.5+c
4*a+5*c=33
4*4.5+5*3=33
18.0+15=33
33.0=33
ok
codetickvd